(xy)^x=e, solve for y'
1
Expert's answer
2020-10-21T15:10:40-0400

Consider the equation "(xy)^x=e"


Take log of base "e" both sides to obtain,


"ln(xy)^x=ln(e)"


"xln(xy)=1" using "ln(a^n)=nln(a)"


Now differentiate "xln(xy)=1" implicitly with respect to "x" as,


"x\frac{d}{dx}ln(xy)+ln(xy)\frac{d}{dx}(x)=\frac{d}{dx}(1)" using product rule


"x(\frac{y+x\frac{dy}{dx}}{xy})+ln(xy)=0"


Replace "\frac{dy}{dx}=y'" and separate for "y'" as,


"\frac{y+xy'}{y}+ln(xy)=0"


"y+xy'+yln(xy)=0"


"xy'=-y-yln(xy)"


"xy'=-y(1+ln(xy))"


"y'=\frac{-y(1+ln(xy))}{x}"


"y'=-\frac{y(1+ln(xy))}{x}"


Therefore, the first derivative with respect to "x" is "y'=-\frac{y(1+ln(xy))}{x}"